OpenStudy (anonymous):

What is the exact value of the expression 7the square root of 8. + 2the square root of 12. − 3the square root of 2.

OpenStudy (anonymous):

\[7\sqrt{8}+2\sqrt{12}-3\sqrt{2}\] \[7\sqrt{4*2}+2\sqrt{4*3}-3\sqrt{2}\] \[14\sqrt{2}+4\sqrt{3}-3\sqrt{2}\] \[11\sqrt{2}+4\sqrt{3}\]

OpenStudy (amistre64):

it more readable like this: 7 sqrt(8) +2 sqrt(12) −3 sqrt(2) just a thought :)